OpenStudy (spicytaco1):

will fan Convert the fraction 2/5 to a decimal and a percent. Show all of your work and explain your answer.

OpenStudy (idealist10):

2/5=0.4=40% Because 1/5=0.2, to get this in decimal, just do the long division 1 divided by 5, you'll get 0.2, so 2/5=0.4. Or just do the long division 2 divided by 5. You'll get 0.4.